If f0(x) = c (c a constant) for all x, to show f(x) = cx + d (Do not use integration)
1
Expert's answer
2020-05-13T16:43:17-0400

Given f'(x)=c

Consider g(x)=f(x)−cx. Then g'(x)=f'(x)−c=c−c=0.

Therefore g is constant, say g(x)=d.

Then f(x)=cx+d.